Solve the equation: x² + 3x - 54 = 0

I cant do this at all :[

Answer by scott8148(6628)

looking for factors of -54 that sum to 3

like 9 and -6

(x+9)(x-6)=0 __ FOILing this gives the original equation

x+9=0 __ x=-9

x-6=0 __ x=6